# Mutual Coupling in Continuous Aperture Arrays: Physical Modeling and Beamforming Design

### v1

@@ I Introduction
In contrast to SPDAs, the mutual coupling effect in CAPAs has received significantly less attention. In particular, early work examined the spatial degrees of freedom between continuous electromagnetic (EM) volumes and surfaces via eigenfunction analysis [miller2000communicating, dardari2020communicating, 9896943] and investigated corresponding capacity limits using methods such as Kolmogorov information theory and Fredholm determinant analysis [jensen2008capacity, 8585146, wan2023mutual]. More recently, several studies have addressed beamforming for CAPAs. Owing to the continuous nature of the aperture, CAPA beamforming is formulated with Hilbert-Schmidt operators and functional optimization, which limits the direct reuse of designs developed for SPDAs. To address these challenges, several approaches have been proposed, including wavenumber-domain discretization [9906802, zhang2023pattern], calculus of variations [10910020, 11122426], and subspace expansion [guo2024deep]. However, mutual coupling is not effectively addressed in the aforementioned works. While a recent study provided a wavenumber-domain analysis of CAPA mutual coupling [11006094], it relied on simplifications of omitting both polarization effects and surface dissipation loss. Consequently, a complete CAPA beamforming design that explicitly and rigorously accounts for mutual coupling is still lacking.
@@ II-A Mutual Coupling
Let $\bm{j}_{\mathrm{t}}(\mathbf{s})\in\mathbb{C}^{3\times 1}$ with $\mathbf{s}\in\mathcal{S}$ denote the source current density with wavelength $\lambda$. The corresponding signal frequency is denoted by $f$. The EM power exerted by this source current is given by [11006094]
$P_{\mathrm{em}}=\frac{1}{2}\Re\left\{\int_{\mathcal{S}}\bm{j}^{H}_{\mathrm{t}}(\mathbf{s})\bm{e}_{\mathrm{t}}(\mathbf{s})d\mathbf{s}\right\},$ (3)
where $\bm{e}_{\mathrm{t}}(\mathbf{s})\in\mathbb{C}^{3\times 1}$ is the electric field excited by the source current across the transmit surface. Specifically, this field consists of two components, namely the radiated field and the dissipated field.
II-A1 Coupling Kernel
Both the radiated and dissipated fields are relevant to the source current density $\bm{j}_{\mathrm{t}}(\mathbf{s})$. Specifically, the radiated field can be characterized by solving the inhomogeneous Helmholtz wave equation, leading to the solution given by [dardari2020communicating, bjornson2024towards, 11006094]
@@ II-A1 Coupling Kernel
where $\mathbf{G}(\mathbf{s})\in\mathbb{C}^{3\times 3}$ is the tensor Green’s function defined by
@@
On the contrary, the dissipative field is introduced by the inherent surface resistance $Z_{s}\in\mathbb{R}$, which is given by
$\bm{e}_{\mathrm{diss}}(\mathbf{s})=Z_{s}\bm{j}_{\mathrm{t}}(\mathbf{s}).$ (7)
Under the assumption that the transmit surface is a good conductor, the surface resistance can be expressed as [pozar2021microwave, Equations (1.125) & (1.131)]
@@
Then, the field $\bm{e}_{\mathrm{t}}(\mathbf{s})$ is given by
$\bm{e}_{\mathrm{t}}(\mathbf{s})$ $=\bm{e}_{\mathrm{diss}}(\mathbf{s})+\bm{e}_{\mathrm{rad}}(\mathbf{s})$
$=\int_{\mathcal{S}}\mathbf{C}(\mathbf{s}-\mathbf{z})\bm{j}_{\mathrm{t}}(\mathbf{z})d\mathbf{z}.$ (9)
The coupling kernel, $\mathbf{C}(\mathbf{s})\in\mathbb{C}^{3\times 3}$, can be expressed as
@@
where $\delta(\mathbf{s})$ is the three-dimensional (3D) Dirac delta function. Substituting (II-A1) into (3) yields the EM transmit power
$P_{\mathrm{em}}$ $=\frac{1}{2}\Re\left\{\int_{\mathcal{S}}\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}^{H}(\mathbf{s})\mathbf{C}(\mathbf{s}-\mathbf{z})\bm{j}_{\mathrm{t}}(\mathbf{z})d\mathbf{z}d\mathbf{s}\right\}$
$=\frac{1}{2}\int_{\mathcal{S}}\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}^{H}(\mathbf{s})\Re\left\{\mathbf{C}(\mathbf{s}-\mathbf{z})\right\}\bm{j}_{\mathrm{t}}(\mathbf{z})d\mathbf{z}d\mathbf{s},$ (11)
where the last step is obtained following [11006094, Appendix B].
@@ II-B Uni-Polarized Coupling Kernel
$\cos^{2}\psi\left((\epsilon^{2}-3)\sin\epsilon+3\epsilon\cos\epsilon\right)+2(\sin\epsilon-\epsilon\sin\epsilon)=0.$ (17)
@@
Based on Proposition 1, we can analyze the nulls of $c_{\mathrm{rad}}(\mathbf{s})$ along the primary axes. The kernel nulls along the $x$-axis are found by setting $\psi=0$ ($i.e.$, $s_{y}=0$) in (17), which yields
@@
These results confirm the simulation in Fig. 2, showing that the nulls are only slightly shifted from the integer half-wavelengths ($i.e.$, $0.5,1.0,1.5,\dots$) of the isotropic case. Similarly, the nulls along the $y$-axis are found by setting $\psi=\pi/2$ ($i.e.$, $s_{x}=0$) in (17), yielding
@@ III-A Optimal Beamforming
The objective of beamforming design is to optimize $w(\mathbf{s})$ to maximize the signal strength at a target receiver, subject to a constraint on the total transmit power $P_{\mathrm{t}}$. This can be formulated as the functional optimization problem given by
@@ III-B1 Wavenumber-domain Approximation
This expression is derived using the Weyl identity [chew1999waves] and the property that the spatial derivative operator $\partial^{2}_{y}$ corresponds to multiplication by $\kappa_{y}^{2}$ after the Fourier transform. The spatial kernel $c_{\mathrm{rad}}(\mathbf{s})$ is then recovered via the inverse 2D Fourier transform
@@
The proof assembles [10938678, Appendix B]. The details are thus omitted here. ∎
@@ III-C2 Numerical Implementation
IV Array Gain and Beampattern Analysis
IV-A Array Gain Analysis
@@
Please refer to Appendix D. ∎
@@ V-A Mutual Coupling
where $\Delta\mathbf{p}_{nm}=\mathbf{p}_{n}-\mathbf{p}_{m}$ denotes the spacing between the $n$-th and the $m$-th antenna elements. Note that the choice of current profile $a_{\mathrm{t}}(\mathbf{s})$ significantly affects the mutual coupling, and it is possible to design $a_{\mathrm{t}}(\mathbf{s})$ carefully to minimize the mutual coupling between different elements.
@@ Appendix A Radiation Power via Poynting Vector
where $\hat{\bm{\kappa}}=\kappa_{0}\mathbf{k}(\theta,\phi)$, similar to (24). Plugging (A) into (93) gives
@@ Appendix B Proof of Proposition 1
$\partial^{2}_{y}\varphi(\mathbf{s})=\frac{d^{2}\varphi}{dr^{2}}\sin^{2}\psi+\frac{d\varphi}{dr}\frac{\cos^{2}\psi}{r},$ (98)
@@
$\cos^{2}\psi\left((\epsilon^{2}-3)\sin\epsilon+3\epsilon\cos\epsilon\right)+2(\sin\epsilon-\epsilon\sin\epsilon)=0.$ (101)
@@ Appendix C Proof of Theorem 1
Appendix D Proof of Proposition 4
@@
Plugging (D) into (108) yields

### v2

# arXiv:2511.11225v2[cs.IT] 27 Feb 2026

@@ I Introduction
In contrast to SPDAs, the mutual coupling effect in CAPAs has received significantly less attention. In particular, early work examined the spatial degrees of freedom between continuous electromagnetic (EM) volumes and surfaces via eigenfunction analysis [miller2000communicating, dardari2020communicating, 9896943] and investigated corresponding capacity limits using methods such as Kolmogorov information theory and Fredholm determinant analysis [jensen2008capacity, 8585146, wan2023mutual]. More recently, several studies have addressed beamforming for CAPAs. Owing to the continuous nature of the aperture, CAPA beamforming is formulated with Hilbert-Schmidt operators and functional optimization, which limits the direct reuse of designs developed for SPDAs. To address these challenges, several approaches have been proposed, including wavenumber-domain discretization based on Fourier plane-wave expansion [9906802, zhang2023pattern], calculus of variations leading to Fredholm integral equations [10910020, 11122426], and subspace expansion [guo2024deep]. However, mutual coupling is not effectively addressed in the aforementioned works. While a recent study provided a wavenumber-domain analysis of CAPA mutual coupling [11006094], it relied on simplifications of omitting both polarization effects and surface dissipation loss, and it did not provide an explicit beamforming solution. Consequently, a complete CAPA beamforming design that explicitly and rigorously accounts for mutual coupling is still lacking.
@@ II-A Mutual Coupling
Let $\bm{j}_{\mathrm{t}}(\mathbf{s})\in\mathbb{C}^{3\times 1},\mathbf{s}\in\mathcal{S}$, denote the Fourier transform of the source current density at frequency $f$, where the corresponding wavelength is $\lambda$. The time factor is assumed to be $e^{-2\pi ft}$. We focus on a narrowband single-carrier system. Wideband multi-carrier systems can be handled by applying the proposed model to each carrier frequency independently.
The total EM power, denoted by $P_{\mathrm{em}}$, exerted by the source current is divided into two fundamental components: the power radiated into free space and the power dissipated as heat due to the finite conductivity of the surface, as discussed below.
II-A1 Radiated Power
The free-space radiated field caused by the source current $\bm{j}_{\mathrm{t}}(\mathbf{s})$ can be characterized by solving the inhomogeneous Helmholtz wave equation, leading to the solution given by [dardari2020communicating, bjornson2024towards, 11006094]
@@ II-A1 Radiated Power
where $\mathbf{G}(\mathbf{s})\in\mathbb{C}^{3\times 3}$ is the dyadic Green’s function:
@@
The radiated power is the time-averaged work done by the source current working against this radiated field across the transmit surface [11006094]:
$P_{\mathrm{rad}}$ $=\frac{1}{2}\Re\left\{\int_{\mathcal{S}}\bm{j}^{H}_{\mathrm{t}}(\mathbf{s})\bm{e}_{\mathrm{rad}}(\mathbf{s})d\mathbf{s}\right\}$
$=\frac{1}{2}\Re\left\{\int_{\mathcal{S}}\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}^{H}(\mathbf{s})\mathbf{G}(\mathbf{s}-\mathbf{z})\bm{j}_{\mathrm{t}}(\mathbf{z})d\mathbf{z}d\mathbf{s}\right\}$
$=\frac{1}{2}\int_{\mathcal{S}}\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}^{H}(\mathbf{s})\Re\left\{\mathbf{G}(\mathbf{s}-\mathbf{z})\right\}\bm{j}_{\mathrm{t}}(\mathbf{z})d\mathbf{z}d\mathbf{s},$ (6)
where the last step is obtained following [11006094, Appendix B].
II-A2 Surface Dissipation Loss
Practical transmit surfaces are not perfect conductors. They exhibit an inherent surface resistance $Z_{s}\in\mathbb{R}$. Under the assumption that the transmit surface is a good conductor, the surface resistance can be expressed as [pozar2021microwave, Eq. (1.125)]
@@ II-A2 Surface Dissipation Loss
Due to this resistance, a portion of the supplied energy is lost strictly to local Joule heating, leading to the following dissipated power [pozar2021microwave, Eq. (1.131)]:
$P_{\mathrm{diss}}=\frac{Z_{s}}{2}\int_{\mathcal{S}}\|\bm{j}_{\mathrm{t}}(\mathbf{s})\|^{2}d\mathbf{s}.$ (8)
II-A3 Coupling Kernel
The total EM power is the sum of both components, calculated by $P_{\mathrm{em}}=P_{\mathrm{rad}}+P_{\mathrm{diss}}$. By factoring out the source current density, we can define a unified coupling kernel $\mathbf{C}(\mathbf{s})\in\mathbb{C}^{3\times 3}$ as
@@ II-A3 Coupling Kernel
where $\delta(\mathbf{s})$ is the three-dimensional (3D) Dirac delta function.
Utilizing this unified coupling kernel, the total EM power can be expressed in a single integral equation as follows:
$P_{\mathrm{em}}=\frac{1}{2}\int_{\mathcal{S}}\int_{\mathcal{S}}\bm{j}_{\mathrm{t}}^{H}(\mathbf{s})\Re\left\{\mathbf{C}(\mathbf{s}-\mathbf{z})\right\}\bm{j}_{\mathrm{t}}(\mathbf{z})d\mathbf{z}d\mathbf{s}.$ (10)
@@ II-B Uni-Polarized Coupling Kernel
$\sin^{2}\psi\left((\epsilon^{2}-3)\sin\epsilon+3\epsilon\cos\epsilon\right)+2(\sin\epsilon-\epsilon\cos\epsilon)=0.$ (16)
@@
Based on Proposition 1, we can analyze the nulls of $c_{\mathrm{rad}}(\mathbf{s})$ along the primary axes. The kernel nulls along the $x$-axis are found by setting $\psi=\pi/2$ ($i.e.$, $s_{y}=0$) in (16), which yields
@@
These results confirm the simulation in Fig. 2, showing that the nulls are only slightly shifted from the integer half-wavelengths ($i.e.$, $0.5,1.0,1.5,\dots$) of the isotropic case. Similarly, the nulls along the $y$-axis are found by setting $\psi=0$ ($i.e.$, $s_{x}=0$) in (16), yielding
@@ III-A Optimal Beamforming
The objective of beamforming design is to optimize $w(\mathbf{s})$ to maximize the signal strength at a target receiver, and consequently, the fundamental single-user capacity limit, subject to a constraint on the total transmit power $P_{\mathrm{t}}$. This can be formulated as the following functional optimization problem:
@@ III-B1 Wavenumber-domain Approximation
This expression is derived using the Weyl identity [chew1999waves] and the property that the spatial derivative operator $\partial^{2}_{y}$ corresponds to multiplication by $\kappa_{y}^{2}$ after the Fourier transform. The detailed derivation of (32) is given in Appendix The spatial kernel $c_{\mathrm{rad}}(\mathbf{s})$ is then recovered via the inverse 2D Fourier transform
@@
The approximated kernel (III-B1) admits the invertible structure described in [10938678, Lemma 2]. Consequently, the closed-form inverse provided therein can be directly applied to obtain (41), which completes the proof. ∎
@@ III-C2 Numerical Implementation
Remark 4.
(Applicability to Near-Field Channels) While this work primarily evaluates beamforming under the far-field channel approximation in (24), the proposed optimization frameworks can be discussed in the context of near-field communications. Specifically, the CG method is inherently channel-agnostic and can be directly applied to the exact near-field spherical wave channel model formulated in (II-C). This requires no algorithmic modifications. Conversely, the KA method leverages the far-field plane-wave assumption to analytically evaluate the spatial integrals as a product of $\mathrm{sinc}$ functions, yielding the closed-form solutions in (2) and (III-B2). If applied to the near-field, this specific spatial integral would lose its closed-form tractability, necessitating numerical integration, such as Gauss-Legendre quadrature, to compute the optimal continuous beamformer.
IV Array Directivity and Beampattern Analysis
IV-A Array Directivity Analysis
@@
Please refer to Appendix E. ∎
@@ V-A Mutual Coupling
where $\Delta\mathbf{p}_{nm}=\mathbf{p}_{n}-\mathbf{p}_{m}$ denotes the spacing between the $n$-th and the $m$-th antenna elements. Note that the choice of current profile $a_{\mathrm{t}}(\mathbf{s})$ significantly affects the mutual coupling, and it is possible to design $a_{\mathrm{t}}(\mathbf{s})$ carefully to minimize the mutual coupling between different elements. Moreover, the model in (83) is consistent with the mutual-coupling impedance matrices commonly used in circuit-based analyses of SPDAs, e.g., [10547020, Eq. (5)].
@@ VII Conclusions
VIII Acknowledgment
The authors would like to thank Dr. Chongjun Ouyang from Queen Mary University of London for the fruitful discussions regarding the initial modeling of mutual coupling.
@@ Appendix A Radiation Power via Poynting Vector
where $\hat{\bm{\kappa}}=\kappa_{0}\mathbf{k}(\theta,\phi)$, similar to (23). Plugging (A) into (92) and using the identity $\big(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}\big)^{2}=\big(\mathbf{I}_{3}-\frac{1}{\kappa_{0}^{2}}\hat{\bm{\kappa}}\hat{\bm{\kappa}}^{T}\big)$ gives
@@ Appendix B Proof of Proposition 1
$\partial^{2}_{y}\varphi(\mathbf{s})=\frac{d^{2}\varphi}{dr^{2}}\cos^{2}\psi+\frac{d\varphi}{dr}\frac{\sin^{2}\psi}{r},$ (97)
@@
$\sin^{2}\psi\left((\epsilon^{2}-3)\sin\epsilon+3\epsilon\cos\epsilon\right)+2(\sin\epsilon-\epsilon\cos\epsilon)=0.$ (100)
@@ Appendix C Proof of Theorem 1
Appendix D Derivation of (32)
This appendix derives the closed-form wavenumber-domain representation in (32) for the radiation mutual-coupling kernel. Define $\Phi(\bm{\kappa})\triangleq\mathcal{F}\{\varphi\}(\bm{\kappa})$. Then, from (15) and the Fourier-domain identity $\mathcal{F}\{\partial_{y}^{2}\varphi\}(\bm{\kappa})=-\kappa_{y}^{2}\Phi(\bm{\kappa})$, the spectrum $C_{\mathrm{rad}}(\bm{\kappa})$ satisfies
$C_{\mathrm{rad}}(\bm{\kappa})=$ $\ \kappa_{0}Z_{0}\left(\Phi(\bm{\kappa})+\frac{1}{\kappa_{0}^{2}}\mathcal{F}\{\partial_{y}^{2}\varphi\}(\bm{\kappa})\right)$
$=$ $\ \kappa_{0}Z_{0}\left(1-\frac{\kappa_{y}^{2}}{\kappa_{0}^{2}}\right)\Phi(\bm{\kappa}).$ (106)
Therefore, it suffices to compute $\Phi(\bm{\kappa})$.
Noting that $\varphi(\mathbf{s})=\Im\{g(\mathbf{s})\}$, we first invoke the Weyl identity at $s_{z}=0$ to express $g(\mathbf{s})$ as [chew1999waves]
$g(\mathbf{s})=\frac{j}{8\pi^{2}}\iint_{-\infty}^{+\infty}\frac{e^{j\bm{\kappa}^{T}\mathbf{s}}}{k_{z}(\bm{\kappa})}\,d\bm{\kappa},\quad k_{z}(\bm{\kappa})=\sqrt{\kappa_{0}^{2}-\|\bm{\kappa}\|^{2}}.$ (107)
For $\|\bm{\kappa}\|\leq\kappa_{0}$, $k_{z}(\bm{\kappa})$ is real. For $\|\bm{\kappa}\|>\kappa_{0}$, $k_{z}(\bm{\kappa})=j\sqrt{\|\bm{\kappa}\|^{2}-\kappa_{0}^{2}}$ is purely imaginary. In the latter case, $j/k_{z}(\bm{\kappa})$ is purely real and hence does not contribute to $\Im\{g(\mathbf{s})\}$. Consequently, $\varphi(\mathbf{s})$ is supported only on the disk $\|\bm{\kappa}\|\leq\kappa_{0}$, which yields
$\varphi(\mathbf{s})=\frac{1}{(2\pi)^{2}}\iint_{\|\bm{\kappa}\|\leq\kappa_{0}}\frac{e^{j\bm{\kappa}^{T}\mathbf{s}}}{2\sqrt{\kappa_{0}^{2}-\|\bm{\kappa}\|^{2}}}\,d\bm{\kappa}.$ (108)
Comparing (108) with the inverse Fourier transform gives
$\Phi(\bm{\kappa})=\begin{cases}\dfrac{1}{2\sqrt{\kappa_{0}^{2}-\|\bm{\kappa}\|^{2}}},&\|\bm{\kappa}\|\leq\kappa_{0},\\[6.0pt] 0,&\|\bm{\kappa}\|>\kappa_{0}.\end{cases}$ (109)
Substituting (109) into (D) yields
$C_{\mathrm{rad}}(\bm{\kappa})=\frac{Z_{0}\left(1-\kappa_{y}^{2}/\kappa_{0}^{2}\right)}{2\sqrt{1-\|\bm{\kappa}\|^{2}/\kappa_{0}^{2}}},\quad\|\bm{\kappa}\|\leq\kappa_{0},$ (110)
and $C_{\mathrm{rad}}(\bm{\kappa})=0$ for $\|\bm{\kappa}\|>\kappa_{0}$. This completes the derivation of (32).
Appendix E Proof of Proposition 4
@@
Plugging (E) into (112) yields